# The Coordination Gap: Alternation Metrics for Temporal Dynamics in Multi-Agent Battle of the Exes

### 10 Mar 2026 v2

@@ 5.5.3 Perfect Alternation Equivalent Analysis
Table 5: Perfect Alternation Equivalent Analysis (CALT - BASE=1000)
Agents Reward CALT ALT Ratio PA Equiv. Performance
Type (QL) Agents (% of Perfect)
2 ILF 0.322 0.568 1.14 56.8%
2 IQF 0.322 0.568 1.14 56.8%
3 ILF 0.142 0.377 1.13 37.7%
3 IQF 0.142 0.377 1.13 37.7%
5 ILF 0.062 0.250 1.25 25.0%
5 IQF 0.062 0.250 1.25 25.0%
8 ILF 0.047 0.216 1.73 21.6%
8 IQF 0.047 0.216 1.73 21.6%
10 ILF 0.048 0.219 2.19 21.9%
10 IQF 0.048 0.219 2.19 21.9%
Interpretation: The ALT Ratio maps observed CALT values to equivalent perfectly alternating agents. For example, 10 Q-learning agents achieve CALT = 0.048, equivalent to only 2.19 perfectly alternating agents (21.9% of perfect coordination).
PA-Equivalent Formula (CALT): $\text{ALT Ratio}=\sqrt{\text{CALT}-1.879\times 10^{-10}}$
$\text{PA Equivalent Agents}=n\times\text{ALT Ratio}$
Key Trend: Performance drops sharply from 2 to 5 agents and remains low for 8–10 agents, demonstrating coordination difficulty scales dramatically with system size.
@@ 6.2 Coordination Breakdown as Strategic Regime Shift
The performance drop from 56.8% (2 agents) to $\sim$22% (8–10 agents) reflects a fundamental shift in the equilibrium structure of multi-agent games: as agent count increases, the strategic complexity of achieving coordinated alternation grows sharply, and independent learners fail to discover equilibria supporting fair turn-taking (Pikovsky, 2009; Maruta and Okada, 2012; Masiliūnas, 2019; Zhu et al., 2020; Shan et al., 2019; He et al., 2015; Tian et al., 2026). Below a certain agent threshold, residual coordination persists through partial turn-taking patterns; beyond this threshold, individual rationality leads to collective irrationality, as agents interfere with each other’s strategies, producing coordination outcomes qualitatively close to random dynamics.
@@ 7.3 Implications, Limitations, and Future Directions
Our results highlight constraints of independent tabular learning for temporal coordination and point to promising directions, including richer state representations, explicit communication mechanisms, and centralized training with decentralized execution. Extended implications, limitations, future directions, and application contexts are provided in Supplementary Section S6. We repeated runs with different random seeds and observed qualitatively consistent patterns; for brevity, we report representative results.
By establishing rigorous baselines, developing interpretable benchmarking methodologies, and exposing limitations in conventional evaluation practices, this work advances both theoretical understanding and practical assessment of multi-agent coordination. More broadly, our findings contribute to the social choice and welfare measurement literature by demonstrating that distributional fairness criteria—widely used in fair division and cooperative game theory—can be fundamentally misleading in repeated interaction settings where the temporal sequence of allocations matters as much as their aggregate distribution. The result that Q-learning agents perform below random baselines (despite appearing successful by traditional measures) underscores the necessity of domain-appropriate metrics and the dangers of relying on outcome fairness alone when procedural fairness matters.

### 21 Mar 2026 v4

@@ 5.5.3 Perfect Alternation Equivalent Analysis
Table 5: Perfect Alternation Equivalent Analysis (CALT, BASE=1000; Type-B state representation). $\text{ALT Ratio}=\sqrt{\text{CALT}-1.879\times 10^{-10}}$; $\text{PA Equiv.}=n\times\text{ALT Ratio}$. IQF (tie reward $r_{\mathrm{high}}/n^{2}$) consistently yields slightly higher CALT than ILF (tie reward $r_{\mathrm{high}}/n$) at intermediate $n$: the lower tie incentive under IQF exerts marginally stronger pressure toward exclusive wins, improving coordination structure. However, the effect is small ($\Delta\leq 0.004$) and both reward structures produce a sharp degradation from $\sim$57% at $n=2$ to $\sim$21% at $n\geq 8$, confirming that coordination failure is reward-structure-agnostic.
Agents Reward CALT (QL) ALT Ratio PA Equiv. Performance
Type Agents (% of Perfect)
2 ILF 0.3222 0.568 1.14 56.8%
IQF 0.3222 0.568 1.14 56.8%
3 ILF 0.1423 0.377 1.13 37.7%
IQF 0.1459 0.382 1.15 38.2%
5 ILF 0.0624 0.250 1.25 25.0%
IQF 0.0631 0.251 1.26 25.1%
8 ILF 0.0467 0.216 1.73 21.6%
IQF 0.0471 0.217 1.74 21.7%
10 ILF 0.0482 0.219 2.19 21.9%
IQF 0.0481 0.219 2.19 21.9%
@@ 6.2 Coordination Breakdown as Strategic Regime Shift
The performance drop from 56.8% (2 agents) to $\sim$22% (8–10 agents) reflects a fundamental shift in the equilibrium structure of multi-agent games: as agent count increases, the strategic complexity of achieving coordinated alternation grows sharply, and independent learners fail to discover equilibria supporting fair turn-taking (Pikovsky, 2009; Maruta and Okada, 2012; Masiliūnas, 2019). Below a certain agent threshold, residual coordination persists through partial turn-taking patterns; beyond this threshold, individual rationality leads to collective irrationality, as agents interfere with each other’s strategies, producing coordination outcomes qualitatively close to random dynamics.
@@ 7.3 Implications, Limitations, and Future Directions
Our results highlight constraints of independent tabular learning for temporal coordination and point to promising directions, including richer state representations, explicit communication mechanisms, and centralized training with decentralized execution. Extended implications, limitations, future directions, and application contexts are provided in Supplementary Section S6. Results are qualitatively consistent across all four configurations (Type-A/B $\times$ ILF/IQF); coordination failure is not an artifact of any specific state representation or reward structure.
By establishing rigorous baselines, developing interpretable benchmarking methodologies, and exposing limitations in conventional evaluation practices, this work advances both theoretical understanding and practical assessment of multi-agent coordination. More broadly, our findings contribute to the social choice and welfare measurement literature by demonstrating that distributional fairness criteria, widely used in fair division and cooperative game theory, can be fundamentally misleading in repeated interaction settings where the temporal sequence of allocations matters as much as their aggregate distribution. The result that Q-learning agents perform below random baselines (despite appearing successful by traditional measures) underscores the necessity of domain-appropriate metrics and the dangers of relying on outcome fairness alone when procedural fairness matters.
A convergent body of work across computational social choice, voting theory, and decision-making has established that temporal structure must be explicitly incorporated into fairness evaluation: temporal fair division (Cookson et al., 2025; Elkind et al., 2025; Choi and Li, 2026), picking sequence analysis (Bouveret and Lang, 2014; Kalinowski et al., 2013), perpetual voting (Lackner, 2020), and temporally aware optimization (Torres et al., 2024) all demonstrate that cumulative criteria alone are insufficient. Our results complement this line of work from a diagnostic perspective: rather than designing mechanisms to enforce temporal fairness, we ask whether self-interested agents can achieve it spontaneously through reinforcement learning, and provide metrics to quantify the gap when they do not. The difficulty of the mechanism design problem itself motivates this diagnostic approach: Elkind et al. (2025) prove that determining whether a TEF1 allocation exists is NP-hard in general, and Choi and Li (2026) show that strict temporal fairness remains impossible in most settings even when a scheduling buffer is available, results that underscore why empirical emergence cannot be assumed and must instead be measured. Choi and Li (2026) further identify strategy-proofness under competitive settings as an open question; our results provide an empirical answer for the reinforcement learning analogue: independent Q-learning agents systematically fail to achieve temporal fairness, with CALT and AALT declining 57–80% relative to random baselines across agent counts in our experimental setting. For comparison, Bouveret and Lang (2014) establish that even deliberate worst-case strategic manipulation of an alternating picking sequence degrades social welfare by at most $\sim$33%; that spontaneous coordination failure in our setting exceeds this bound illustrates a broader point: without temporally sensitive metrics, failures of this magnitude would remain entirely invisible, misclassified as successful coordination by traditional measures.